OpenStudy (anonymous):

Which of the following rational numbers is equal to 4.7 (repeated 7)? A). 43/9 B). 49/9 C). 35/9 D). 32/9 THIS IS IMPORTANT I NEED AN ANSWER

OpenStudy (anonymous):

to find you just need to do the following 1) let x=4.7 (7 repeated) 2) then 10x=4.7(7)*10=47.7(7) Note: by notation (7) i mean 7 repeated, but it isn't the notation that is used in US. In US there is used line over the number. 3) then solve for 10x-x 4) 10x-x = 47.(7) - 4.(7) 9x=43 so now find x

OpenStudy (anonymous):

ya but it wouldnt let me put a line over it & do you know what the answer is ?

OpenStudy (anonymous):

the asnwer is 43/9.

OpenStudy (anonymous):

i think its 43/9

OpenStudy (anonymous):

ok thanksss

OpenStudy (anonymous):

you can check it finally by calc if u want to.

OpenStudy (anonymous):

thank you so much!